OpenStudy (anonymous):

if f(x) = 3x + 5 and g(x) = 4 - 5x, determine the values of x such that f(g(x)) = g(f(x)) if f(x) = 3x + 5 and g(x) = 4 - 5x, determine the values of x such that f(g(x)) = g(f(x)) @Mathematics

OpenStudy (anonymous):

f(g(x))=3*(4-5x)+5 g(f(x))=4-5*(3x+5) 3*(4-5x)+5=4-5*(3x+5) 12-15*x+5=4-15*x-25 17-15*x=-21-15*x 38=0x?!

OpenStudy (anonymous):

I'm not getting it either!

OpenStudy (anonymous):

the answer is that there is no value of x that accomplish that equality that's it in my opinion

OpenStudy (anonymous):

Alright